​(Calculating free cash flows​) At​ present, Solartech Skateboards is considering expanding its product line to include​ gas-powered skateboards;​ however, it is questionable how well they will be received by skateboarders. Although you feel there is a 50 percent chance you will sell 11 comma 000 of these per year for 10 years​ (after which time this project is expected to shut down because​ solar-powered skateboards will become more​ popular), you also recognize that there is a 25 percent chance that you will only sell 3 comma 000 and also a 25 percent chance you will sell 15 comma 000. The gas skateboards would sell for $ 110 each and have a variable cost of ​$35 each. Regardless of how many you​ sell, the annual fixed costs associated with production would be ​$130 comma 000. In​ addition, there would be an initial expenditure of ​$1 comma 000 comma 000 associated with the purchase of new production equipment. It is assumed that this initial expenditure will be depreciated using the simplified​ straight-line method down to zero over 10 years. Because of the number of stores that will need​ inventory, the working capital requirements are the same regardless of the level of sales. This project will require a​ one-time initial investment of $ 70 comma 000 in net working​ capital, and that​ working-capital investment will be recovered when the project is shut down.​ Finally, assume that the​ firm's marginal tax rate is 32 percent.

a. What is the initial outlay associated with the​ project?

b. What are the annual free cash flows associated with the project for years 1 through 9 under each sales​ forecast? What are the expected annual free cash flows for years 1 through​ 9?

c. What is the terminal cash flow in year 10​ (that is, what is the free cash flow in year 10 plus any additional cash flows associated with the termination of the​ project)?

d. Using the expected free cash​ flows, what is the​ project's NPV given a required rate of return of 12 ​percent? What would the​ project's NPV be if 11 comma 000 skateboards were​ sold?
Answer #1

Part (a)

Initial outlay = purchase of new production equipment + investment in one time working capital = 1,000,000 + 700,000 = $ 1,700,000

Part (b)

Expected sale volume, V = probability weighted average = 50% x 11,000 + 25% x 3,000 + 25% x 15,000 =  10,000

Annual depreciation, D = 1,000,000 / 10 = 100,000

the expected annual free cash flows for years 1 through​ 9 = [(SP - CP) x V - FC - D] x (1 - Tax rate) + D = [(110 - 35) x 10,000 - 130,000 -100,000] x (1 - 32%) + 100,000 = $  453,600

Part (c)

the terminal cash flow in year 10 = Annual free cash flows calculated above + recovery of working capital = 453,600 + 700,000 = $  1,153,600

Part (d)

NPV = - Initial investment + PV of cash flows through year 1 to 9 as annuity + PV of year 10 cash flows = -1,700,000 + A / r x [1 - (1 + r)-n] + 1,153,600 / (1 + r)n + 1 = - 1,700,000 + 453,600 / 12% x [1 - (1 + 12%)-9] + 1,153,000 x (1 + 12%)-10 = - 1,700,000 + 2,416,894 + 371,428 = $  1,088,322

If V = 11,000 then annual cash flows over year 1 to 9 = = [(SP - CP) x V - FC - D] x (1 - Tax rate) + D = [(110 - 35) x 11,000 - 130,000 -100,000] x (1 - 32%) + 100,000 = $ 504,600

Year 10 cash flows =  504,600 + 700,000 = $  1,204,600

Hence, NPV = - Initial investment + PV of cash flows through year 1 to 9 as annuity + PV of year 10 cash flows = -1,700,000 + A / r x [1 - (1 + r)-n] + 1,204,600 / (1 + r)n + 1 = - 1,700,000 + 504,600 / 12% x [1 - (1 + 12%)-9] + 1,204,600 x (1 + 12%)-10 = - 1,700,000 +  2,688,635 +  387,849 = $  1,376,484